OpenStudy (anonymous):

plz help solve for x = x – 3 > –9

OpenStudy (jackellyn):

x-3 >-9 add 3 to both sides x-3+3 >-9+3 ^cancels x >-9+3 What is -9+3?

OpenStudy (anonymous):

-6

OpenStudy (jackellyn):

correct @zahabaha x >-6 understand?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

. –2x + 1 ≤ –11 how wuld u do thtis

OpenStudy (jackellyn):

Okay, I'll go step by step with another example. x+5 > 10 So, you want to get the x alone (that is what solve for x means) To get rid of the 5, you have to do the opposite. The opposite of adding is subtracting. So x+5-5 is going to cancel the 5s and get the x alone and what you do to one side you do to the other sides. x+5-5 > 10-5 x >10-5 x> 5

OpenStudy (anonymous):

ooh now i gwt it

OpenStudy (anonymous):

note when you times or divide by a negative number, remember to switch the side > <

OpenStudy (anonymous):

but the question i asked looks harder

OpenStudy (jackellyn):

For -2x + 1 <= -11 First step: subtract 1 from both sides. -2x +1 -1 <= -11 -1 -2x <= -12 x => 6

OpenStudy (anonymous):

ohhk now i geety it thanks

OpenStudy (jackellyn):

The second step was to divide both sides by -2 which switches the sign.

OpenStudy (anonymous):

wwat dont get tat part so it be's x<6

OpenStudy (jackellyn):

when you divide by a negative number (to get the x alone) you have to switch the sign. examples: -3x > 6 -3x/-3 > 6/-3 x < -2 -5x < -10 -5x/-5 < -10/-5 x > 2

OpenStudy (anonymous):

ok so the sign becomes faceing x =< 6

OpenStudy (jackellyn):

\[x \ge 6\]

OpenStudy (anonymous):

oh ok thanks